If y=(x^n)log(x). Than find nth order derivative of y.

Grade:12th pass

1 Answers

Vikas TU
7 years ago
Find first first order derivative,
dy/dx = x^n/x +nx^(n-1)*logx
d2y/dx2 = (n-1)x^(n-2) + n[1/x * nx^(n-1) + logx * n(n-1)x^(n-2)]
.                        .
.                       .
Genralizing,
y(n) = n!/x somthing like that.
U can also proove it by mathematical Induction.